OpenStudy (anonymous):

Which equation represents the line that contains the point (-2, 9) and is perpendicular to the line y = 2x + 6?

OpenStudy (anonymous):

It should be of the form \[ y=-\frac 1 2 x + b \] Find b so that the line goews through (-2,9)

OpenStudy (anonymous):

\[ 9 = -\frac 1 2 (-2) +b \\ 9=1+b\\ b=8 \]

OpenStudy (anonymous):

Final Answer is \[ y=-\frac 1 2 x +8 \]